# Understanding Higher-Order Correlations Among Semantic Components in Embeddings

### v1

@@ Abstract
Independent Component Analysis (ICA) is an effective method for interpreting the intrinsic geometric structure of embeddings as semantic components. While ICA theory assumes that embeddings can be linearly decomposed into independent components, real-world data often do not satisfy this assumption. Consequently, there are remaining non-independencies between the estimated components that ICA cannot eliminate. We quantified these non-independencies using higher-order correlations and demonstrated that when the higher-order correlation between two components is large, it indicates a strong semantic association between them. The entire structure was revealed through visualization using a maximum spanning tree of semantic components. These findings allow for further understanding of embeddings through ICA.
@@ Results.
Table 3 presents component pairs selected from the maximum spanning tree $T_{150}$ (Sec. 5) and words significantly contributing to their $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values. These words often relate to the meanings of both components. For example, in the Axis 10 and Axis 2 pair, words like ribose, deoxyribose, phosphodiester, biosynthesis, methyltransferase, and pyrimidine notably contribute to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ value, linking biomolecules and chemical components. Detailed results are shown in Appendix G.
@@
Figure 4 shows word embedding scatter plots for axis pairs (10, 2) and (27, 64) with large higher-order correlations to illustrate the distribution of words with significant contributions to higher-order correlations. Unlike a typical independent component scatter plot (Fig. 2), these exhibit many words with large component values in both axes. For the (10, 2) pair, words that notably contribute to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ (ribose, deoxyribose, phosphodiester, biosynthesis, methyltransferase, and pyrimidine) appear with significant values in both components. This abundance of words sharing both semantic components is characteristic of pairs with large higher-order correlations. Detailed results are shown in Appendix C.
@@ C.2 Scatterplots for Independent Axes
Appendix D Prompt Used for Evaluation by GPT

### v2

@@ Abstract
Independent Component Analysis (ICA) offers interpretable semantic components of embeddings. While ICA theory assumes that embeddings can be linearly decomposed into independent components, real-world data often do not satisfy this assumption. Consequently, non-independencies remain between the estimated components, which ICA cannot eliminate. We quantified these non-independencies using higher-order correlations and demonstrated that when the higher-order correlation between two components is large, it indicates a strong semantic association between them, along with many words sharing common meanings with both components. The entire structure of non-independencies was visualized using a maximum spanning tree of semantic components. These findings provide deeper insights into embeddings through ICA.
@@ Results.
Table 3 presents component pairs selected from the maximum spanning tree $T_{150}$ (Sec. 5) and words significantly contributing to their $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ values. These words often relate to the meanings of both components, demonstrating additive compositionality. For example, in the Axis 10 and Axis 2 pair, words like ribose, deoxyribose, phosphodiester, biosynthesis, methyltransferase, and pyrimidine notably contribute to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ value, linking biomolecules and chemical components. Detailed results are shown in Appendix G.
@@
Figure 4 shows word embedding scatter plots for axis pairs (10, 2) and (27, 64) with large higher-order correlations to illustrate the distribution of words with significant contributions to higher-order correlations. Unlike a typical independent component scatter plot (Fig. 2), these exhibit many words with large component values in both axes, reflecting the meanings of both axes and demonstrating the additive compositionality of embeddings. For the (10, 2) pair, words that notably contribute to the $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ (ribose, deoxyribose, phosphodiester, biosynthesis, methyltransferase, and pyrimidine) appear with significant values in both components. This abundance of words sharing both semantic components is characteristic of pairs with large higher-order correlations. Detailed results are shown in Appendix C.
@@ C.2 Scatterplots for Independent Axes
Words with Significant Contributions to Higher-Order Correlations.
We have observed words with significant contributions, i.e., with large values of $\mathbf{S}_{t,i}^{2}\mathbf{S}_{t,j}^{2}$, to higher-order correlations $\mathrm{E}(S_{i}^{2}S_{j}^{2})$ in Table 3 in Sec. 4.3, and will see further examples in Tables 8 and 9 in Appendix G. Such words are labeled in red in the scatter plots in Fig. 4 in Section 4.3, as well as in Figs. 10 and 11 in Appendix C.2. For axis pairs with large higher-order correlations, we observe a large number of words that make significant contributions to the higher-order correlations. The meanings of these words include both axes’ meanings, demonstrating the additive compositionality of embeddings.
Appendix D Prompt Used for Evaluation by GPT-4o mini
@@
When conducting the experiment, we took the following steps to eliminate any potential biases arising from the order of word sequences and specific label names.
  1. 1.
    To remove the influence of word order within lists, we randomly shuffled the words in wordlist_1, wordlist_2, and wordlist_3.
  2. 2.
    To prevent bias in output labels, we conducted the experiment twice, swapping wordlist_2 and wordlist_3 between runs.
  3. 3.
    To account for order bias in the prompt, we randomly alternated the order of the following two lines:
    List pair (A, B): ({wordlist_1}, {wordlist_2})
    List pair (C, D): ({wordlist_1}, {wordlist_3})